OpenStudy (anonymous):

Which quadratic equation has a vertex of (0, 5)?

OpenStudy (anonymous):

x^2-2x+1=0

OpenStudy (anonymous):

Well, there are quite a few that might qualify, but \[y=ax^2+5\]with \[a \in \mathbb{R}\]seems to be a general statement of most of the suspects.

OpenStudy (anonymous):

Further note: \[a \in \mathbb{R}/\left\{ 0 \right\}\]

OpenStudy (anonymous):

Vertex form is y = (x - h)^2 + k where (h, k) is the vertex so for vertex (0, 5) y = (x - 0)^2 + 5 y = x^2 + 5